• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Eugene

    Log entry time 19:44:28 on January31,2002

    Entry number 81398

    keyword=RCS calorimeter repair, HV changed

    At about 16:00 a LG channel #286 tripped. The problem was with the voltage
    divider. It was repared by about 18:30 (by Albert, Igor, Yuri etc).
    At the same time a filter card was fixed, which lead to a changed
    in HV for LG#135,137, using the laser signals.

    New HV settings are stored in the database (marked by the time)
    and to a file "hv_setting_04".